OpenStudy (anonymous):

Solve: ln 1/2 x = 1/4

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

lol woops didnt see ln

OpenStudy (anonymous):

e ^1/4 = 1/2 x 2 (e ^1/4) =x

OpenStudy (anonymous):

thank you!!